Total revenue and cost function of competitive firm are as follows

R=1400Q

C=Q3-60Q2+1500Q+1800

  1. Calculate the amount of profit contribution.
  2. What is break-even level of output?
  3. Determine the output level corresponding to 500 unit profit.
1
Expert's answer
2021-06-17T09:46:43-0400

Solution:

1.). Amount of profit contribution:

Profit contribution = TR – VC

TR = 1400Q

VC = Q3 - 60Q2+1500Q

First determine Q:

Set MR = MC

MR = 1400

MC = 2Q2 - 120Q + 1500

1400 = 2Q2 - 120Q + 1500

Q = 59

Profit contribution = 1400(59) – (593 – 60(592) + 1500(59)

Profit contribution = 82,600 – 85,019 = (2,419)

 

2.). At break-even: TR = TC

1400Q = Q3 - 60Q2+1500Q + 1800

Q = 7 units

 

3.). Output level corresponding to 500-unit profit:

1400Q = Q3 - 60Q2+1500Q + 1800 + 500

Q = 8 Units

 

Output level corresponding to 500-unit profit = 8 units